For the reaction

A(aq) ---> B(aq)

the change in the standard free enthalpy is 2.74 kJ at 25 oC and 5.32 kJ at 45 oC. Calculate the value of this reaction's equilibrium constant at 75 oC.
1
Expert's answer
2020-07-16T06:59:51-0400

ΔG = −RTlnK

K = e−ΔG/RT

K(25) = e2.74/(0.008314×298) = 3.02

K(45) = e5.32/(0.008314×318) = 7.48

The Van't Hoff equation

ln(K2/K1) = (−ΔH/R)(1/T2 − 1/T1)

Enthalpy:

ΔH = −(ln(K2/K1)/(1/T2 − 1/T1))×R = 35.73 kJ/mol

The entropy change for the reaction at 298 K:

ΔG = ΔH − T×ΔS

ΔS = −(ΔG −ΔH)/T = 0.11 kJ/mol×K

Assume that both enthalpy and entropy changes do not change with temperature , and so we can get the Gibbs free energy value at 75 oC:

ΔG = ΔH − T×ΔS = 35.73 − (348×0.11) = −2.55 kJ

The equilibrium constant for this reaction at 75 oC:

K = e−ΔG/RT = e2.55/(0.008314×348) = 2.41

Answer: 2.41